OpenStudy (anonymous):

Which of the following has a slope of –2 and a y-intercept of 4?

OpenStudy (ipwnbunnies):

y = mx + b Where 'm' is the slope of a line, and b is the y-intercept of the line, or where the line crosses the y-axis (vertical axis).

OpenStudy (anonymous):

thanks alot

OpenStudy (mathmale):

Which of the following has a slope of –2 and a y-intercept of 4? This question implies that there are a couple of possible answers / illustrations associated with it. Please post those answers / illustrations.

OpenStudy (anonymous):

A. y = 2x + 4 B. y = –2x – 4 C. y = –2x + 4 D. y = 2x – 4